OpenStudy (anonymous):

Use the definitions sinθ=y/r, cosθ=x/r and/or tanθ=y/x to prove that cotθsecθ=cscθ.

OpenStudy (anonymous):

Use the Pythagorean identity sin² θ + cos² θ = 1 to show that 1 + cot² θ = csc ² θ.>> GIVEN sin² θ + cos² θ = 1 Divide both sides by sin^2 θ 1 + cos² θ/sin² θ = 1/sin² θ and since cos² θ/sin² θ = cot² θ and 1/sin² θ = csc² θ then the above becomes 1 + cot² θ = csc ² θ **************************************… Use definitions sinθ = y/r, cosθ = x/r and/or tanθ = y/x to prove that cotθsecθ = cscθ NOTE that cotθ = cosθ/sinθ = (x/r)/(y/r) = x/y and secθ = 1/cosθ = 1/(x/r) = r/x and cscθ = 1/sinθ = 1/(y/r) = r/y Then (x/y)( r/x) = r/y Simplifying the left hand side of the equation, r/y = r/y Therefore, since cscθ = r/y, then cotθsecθ = cscθ.